OpenStudy (anonymous):

Simplify:x squared + 2x + 1 over 2x + 2

OpenStudy (anonymous):

x^2 + 2x +1 over 2x +2 x^2 + 2x + 1 = (x+1)(x+1) over 2(x+1) = x+1 over 2